# KOROL: Learning Visualizable Object Feature with Koopman Operator Rollout for Manipulation

### v1

@@ Abstract
Learning dexterous manipulation skills presents significant challenges due to complex nonlinear dynamics that underlie the interactions between objects and multi-fingered hands. Koopman operators have emerged as a robust method for modeling such nonlinear dynamics within a linear framework. However, current methods rely on runtime access to ground-truth (GT) object states, making them unsuitable for vision-based practical applications. Unlike image-to-action policies that implicitly learn visual features for control, we use a dynamics model, specifically the Koopman operator, to learn visually interpretable object features critical for robotic manipulation within a scene. We construct a Koopman operator using object features predicted by a feature extractor and utilize it to auto-regressively advance system states. We train the feature extractor to embed scene information into object features, thereby enabling the accurate propagation of robot trajectories. We evaluate our approach on simulated and real-world robot tasks, with results showing that it outperformed the model-based imitation learning NDP by 1.08$\times$ and the image-to-action Diffusion Policy by 1.16$\times$. The results suggest that our method maintains task success rates with learned features and extends applicability to real-world manipulation without GT object states.
@@ 1 Introduction
Figure 1: left: Vanilla Koopman operators rely on ground-truth state which may be difficult to obtain in real-world settings. right: In contrast, we propose
@@
Humans possess an extraordinary ability to manipulate objects, discerning position, shape, and other properties with just a glance. How can robots be endowed with similar perceptual and dexterous manipulation capabilities? Traditional control and optimization approaches typically require detailed models of the system dynamics [hogan2020tactile, mordatch2012contact]. However, these models can be difficult to derive and often lack the flexibility and generalizability needed to adapt to task or environment changes. End-to-end data-driven methods overcome these challenges by learning actions directly from observations [levine2016end, zhao2023learning, shridhar2023perceiver]. While these methods can make minimal assumptions, they often require a large number of demonstrations to master basic skills due to the high dimensionality of the inputs.
@@
To combine sample efficiency of traditional model-based approaches with high generalizability of deep learning methods, one branch of recent work has focused on learning dynamics models to plan trajectories. These methods embed learning into various models, such as Koopman operator [han2023utility, lusch2018deep], Dynamic Movement Primitives (DMP) [bahl2020neural], Neural Geometric Fabrics [xie2023neural, van2024geometric], and more [nagabandi2018neural], showcasing good performance in simulations. However, they often falter in real-world applications due to their reliance on hard-to-obtain ground-truth (GT) or estimated state information like object poses and contact points. Moreover, the learned dynamics models do not transfer across different tasks without a universal state space design.
We propose an approach to remove the dependency on GT states in model-based manipulation learning (Figure LABEL:fig:teaser). Central to this approach, Koopman Operator Rollout for Object Feature Learning (
@@ 2 Related Work
Imitation learning serves as a primary method for teaching robots to manipulate objects by mapping observations or world states directly to actions. Common approaches include Behavioral Cloning (BC) [pomerleau1988alvinn], Implicit Behavioral Cloning (IBC) [florence2022implicit], Long Short-Term Memory (LSTM) networks [levine2016end, florence2019self], Transformers [zhao2023learning, shridhar2023perceiver], and Diffusion Models [chi2023diffusion]. A significant challenge in imitation learning is representing the visual information of a scene. Strategies include using pre-trained 2D [liu2022instruction] or 3D backbones [shridhar2023perceiver] to output visual embeddings. Other works propose end-to-end learning approaches that simultaneously train the visual encoder and the learning policy [zhao2023learning, chi2023diffusion]. Other techniques focus on learning visual representations through correspondence models [florence2019self], self-supervised novel view reconstruction using Neural Radiance Fields (NeRF) [li20223d, ze2023gnfactor], and Gaussian Splatting [lu2024manigaussian].
@@
In robotics, traditional model-based approaches rely on expert knowledge of physics to design system models [mason1985robot, collins2005efficient, mordatch2012contact]. Since traditional methods can miss complex nonlinearities, and end-to-end learning approaches can be data-intensive, a middle ground of data-driven model learning shows promise as a data-efficient way to derive complex models [han2023utility, deisenroth2013gaussian]. Model learning includes a variety of dynamics models such as Koopman operators [korda2018linear, bevanda2021koopman], Deep Neural Koopman operators [lusch2018deep], Dynamic Movement Primitives [bahl2020neural], Neural Geometric Fabrics [xie2023neural, van2024geometric], and others [nagabandi2018neural, nguyen2009model]. Additionally, some studies focus on learning environmental responses to actions to plan a future trajectory [ke2018modeling, yang2023foundation, plate], integrate planning in a generative modeling process [janner2022planning, chi2023diffusion], and seamlessly blend the learning of models with planning [du2019model, chen2023planning].
@@
In the early 1930s, Koopman and Von Neumann introduced the Koopman operator theory to transform complex, nonlinear dynamics systems into linear ones in an infinite-dimensional vector space, using observables as lifted states [koopman1932dynamical, koopman1931hamiltonian]. This transform allows the application of linear system tools for effective prediction, estimation, and control with hand-designed observables [han2023utility, brunton2016koopman, chang2016compositional, bruder2019nonlinear]. Recent methods using neural networks to learn observables have proven more expressive and effective, particularly in chaotic time-series prediction [lusch2018deep, brunton2021modern, yeung2019learning]. Furthermore, the integration of neural network-derived Koopman observables with Model Predictive Control has shown promise in enhancing control tasks [li2019learning, wang2022koopman]. As a significant benchmark, Han et al. [han2023utility] demonstrate the effectiveness of Koopman operators in manipulation tasks using GT object states. Building on this foundation, we extend the application of the Koopman operator to vision-based manipulation tasks in real-world settings by learning object features directly from images
@@ 3 Background: Koopman Operator Theory
Given a dataset $D$, in which each trajectory $\tau=[\mathrm{x}(1),\mathrm{x}(2),\cdots,\mathrm{x}(T)]$ containing $T$ time steps, we can learn $\mathbf{K}$ by minimizing the state prediction error [brunton2021modern]
@@
Figure 2: Training and Execution Pipeline. During training,
@@ 4 Method
We propose

### v2

# Denotes equal contribution

@@ Abstract
Learning dexterous manipulation skills presents significant challenges due to complex nonlinear dynamics that underlie the interactions between objects and multi-fingered hands. Koopman operators have emerged as a robust method for modeling such nonlinear dynamics within a linear framework. However, current methods rely on runtime access to ground-truth (GT) object states, making them unsuitable for vision-based practical applications. Unlike image-to-action policies that implicitly learn visual features for control, we use a dynamics model, specifically the Koopman operator, to learn visually interpretable object features critical for robotic manipulation within a scene. We construct a Koopman operator using object features predicted by a feature extractor and utilize it to auto-regressively advance system states. We train the feature extractor to embed scene information into object features, thereby enabling the accurate propagation of robot trajectories. We evaluate our approach on simulated and real-world robot tasks, with results showing that it outperformed the model-based imitation learning NDP by 1.08$\times$ and the image-to-action Diffusion Policy by 1.16$\times$. The results suggest that our method maintains task success rates with learned features and extends applicability to real-world manipulation without GT object states. Project video and code are available at: https://github.com/hychen-naza/KOROL.
@@ 1 Introduction
Figure 1: left: Vanilla Koopman operators rely on ground-truth state which may be difficult to obtain in real-world settings. right: In contrast, we propose KOROL, which learns a dynamics model and task-relevant object features without labels of object states. The visualization shows the localization of learned feature around the door handle.
@@
Humans possess an extraordinary ability to manipulate objects, discerning position, shape, and other properties with just a glance. How can robots be endowed with similar perceptual and dexterous manipulation capabilities? Traditional control and optimization approaches typically require detailed models of the system dynamics [1, 2]. However, these models can be difficult to derive and often lack the flexibility and generalizability needed to adapt to task or environment changes. End-to-end data-driven methods overcome these challenges by learning actions directly from observations [3, 4, 5]. While these methods can make minimal assumptions, they often require a large number of demonstrations to master basic skills due to the high dimensionality of the inputs.
@@
To combine sample efficiency of traditional model-based approaches with high generalizability of deep learning methods, one branch of recent work has focused on learning dynamics models to plan trajectories. These methods embed learning into various models, such as Koopman operator [6, 7], Dynamic Movement Primitives (DMP) [8], Neural Geometric Fabrics [9, 10], and more [11], showcasing good performance in simulations. However, they often falter in real-world applications due to their reliance on hard-to-obtain ground-truth (GT) state information like object poses and contact points. Moreover, the problems of using computer vision to estimate these object states are the uncertainty in determining the number of objects to consider and which specific states to estimate. Additionally, the learned dynamics models do not transfer across different tasks without a universal state space design.
We propose an approach to remove the dependency on GT states in model-based manipulation learning (Figure 1). Central to this approach, Koopman Operator Rollout for Object Feature Learning (KOROL), is learning visual features that predict robot states during dynamics model rollouts. Unlike learning methods that learn implicit visual features for image-to-action policies, KOROL explicitly trains on visual object features, encoding essential scene information to enhance predictions of robot states during autoregressive model rollouts. Central to this rollout process is the Koopman operator, which utilizes current object features to advance robot states. This establishes a synergistic relationship between the learned object features and the Koopman operator. KOROL uses trained features to refine the Koopman operator, which in turn improves the feature learning process through a more accurate dynamics model.
In experiments, we demonstrate how KOROL outperforms prior methods in ADROIT Hand [12] and generates interpretable visualizations by learning object features from images. We also show how KOROL enables the application of Koopman operator in vision-based real-world manipulation tasks. Finally, we demonstrate how learning object feature instead of designing object states for dynamics modeling enables KOROL to construct universal Koopman dynamics across multiple manipulation tasks. This paper makes the following contributions:
  • We introduce KOROL, an imitation learning method, that uses the Koopman operator to learn object features and show that the Koopman operator with learned object features can outperform the one with GT object states.
  • We extend the application of the Koopman operator to vision-based manipulation tasks in real-world by learning object features from images and demonstrate its effectiveness through comparisons with prior methods.
  • We demonstrate that KOROL learns dimensionally-aligned object features across tasks, enabling the development of a multi-tasking Koopman operator.
@@ 2 Related Work
Imitation learning serves as a primary method for teaching robots to manipulate objects by mapping observations or world states directly to actions. Common approaches include Behavioral Cloning (BC) [13], Implicit Behavioral Cloning (IBC) [14], Long Short-Term Memory (LSTM) networks [3, 15], Transformers [4, 5], and Diffusion Models [16]. A significant challenge in imitation learning is representing the visual information of a scene. Strategies include using pre-trained 2D [17] or 3D backbones [5] to output visual embeddings. Other works propose end-to-end learning approaches that simultaneously train the visual encoder and the learning policy [4, 16]. Other techniques focus on learning visual representations through correspondence models [15], self-supervised novel view reconstruction using Neural Radiance Fields (NeRF) [18, 19], and Gaussian Splatting [20].
@@
In robotics, traditional model-based approaches rely on expert knowledge of physics to design system models [21, 22, 2]. Since traditional methods can miss complex nonlinearities, and end-to-end learning approaches can be data-intensive, a middle ground of data-driven model learning shows promise as a data-efficient way to derive complex models [6, 23]. Model learning includes a variety of dynamics models such as Koopman operators [24, 25], Deep Neural Koopman operators [7], Dynamic Movement Primitives [8], Neural Geometric Fabrics [9, 10], and others [11, 26]. Additionally, some studies focus on learning environmental responses to actions to plan a future trajectory [27, 28, 29], integrate planning in a generative modeling process [30, 16], and seamlessly blend the learning of models with planning [31, 32].
@@
In the early 1930s, Koopman and Von Neumann introduced the Koopman operator theory to transform complex, nonlinear dynamics systems into linear ones in an infinite-dimensional vector space, using observables as lifted states [33, 34]. This transform allows the application of linear system tools for effective prediction, estimation, and control with hand-designed observables [6, 35, 36, 37]. Recent methods using neural networks to learn observables have proven more expressive and effective, particularly in chaotic time-series prediction [7, 38, 39]. Furthermore, the integration of neural network-derived Koopman observables with Model Predictive Control has shown promise in enhancing control tasks [40, 41]. As a significant benchmark, Han et al. [6] demonstrate the effectiveness of Koopman operators in manipulation tasks using GT object states. Building on this foundation, we extend the application of the Koopman operator to vision-based manipulation tasks in real-world settings by learning object features directly from images
@@ 3 Background: Koopman Operator Theory
Given a dataset $D$, in which each trajectory $\tau=[\mathrm{x}(1),\mathrm{x}(2),\cdots,\mathrm{x}(T)]$ containing $T$ time steps, we can learn $\mathbf{K}$ by minimizing the state prediction error [38]
@@
Figure 2: Training and Execution Pipeline. During training, KOROL updates the feature extractor $f_{\theta}$ based on the loss between the predicted robot trajectory $\hat{\tau}_{r}=[\hat{\mathrm{x}}_{r}(1),\hat{\mathrm{x}}_{r}(2),\cdots,\hat{\mathrm{x}}_{r}(T)]$ obtained through Koopman operator rollouts and the ground-truth robot trajectory $\tau_{r}=[\mathrm{x}_{r}(1),\mathrm{x}_{r}(2),\cdots,\mathrm{x}_{r}(T)]$. KOROL updates the Koopman operator with the new object features $\hat{\mathrm{x}}_{o}(t)$ every $M$ epochs to enhance the training of $f_{\theta}$. During execution, KOROL feeds the generated trajectory to the inverse dynamics controller to produce the actions.
@@ 4 Method
We propose KOROL, which formulates dynamics learning and object feature learning as an imitation (supervised) learning problem on robot states. Given a dataset $D$, in which each trajectory $\tau=[\mathrm{x}_{r}(1),\mathrm{y}(1),\mathrm{x}_{r}(2),\mathrm{y}(2),\cdots,\mathrm{x}_{r}(T),\mathrm{y}(T)]$ containing robot states $\mathrm{x}_{r}(t)$ and image observation $\mathrm{y}(t)$ of the object, instead of object state $\mathrm{x}_{o}(t)$, our goal is to learn a visual object feature extractor $f_{\theta}$ and a Koopman operator $\mathbf{K}$ which can predict object features from images that minimize the robot states imitation errors. In this formulation, (2) becomes
$\operatorname*{arg\,min}_{\theta,\mathbf{K}}$ $\sum_{\mathrm{x}\in D}\sum_{t=0}^{T-1}\left\|\mathrm{x}_{r}(t+1)-\mathbf{K}^{\prime}(\mathrm{x}_{r}(t),f_{\theta}(\mathrm{y}(t)))\right\|^{2}.$ (5)
Learning object feature.
While traditional Koopman operator construction requires GT object state information $\mathrm{x}_{o}(t)$ (section 3), instead, we adopt a neural network $f_{\theta}$ for object feature encoding and extraction from RGBD images. We initialize $f_{\theta}$ and predict object features $\hat{\mathrm{x}}_{o}(t)$ for all images in $D$ to construct the initial Koopman operator $\mathbf{K}$, as proposed by Han et al. [6]. During training, we randomly sample the beginning time step $t_{0}$ in trajectory $\tau$, and select $\hat{\mathrm{x}}_{o}(t_{0})=f_{\theta}(\mathrm{y}(t_{0}))$. Then we can advance the observables forward using $\mathbf{K}\phi(\cdot,\cdot)$. We train $f_{\theta}$ by minimizing the loss function
$\mathcal{L}=$ $\mathbb{E}_{\tau,t_{0}}\left[\sum_{i=0}^{N-1}\left\|\mathrm{x}_{r}(t_{0}+i+1)-\mathbf{K}^{\prime}(\hat{\mathrm{x}}_{r}(t_{0}+i),\hat{\mathrm{x}}_{o}(t_{0}+i)))\right\|^{2}\right],$ (6)
where $N$ is the prediction horizon and $\hat{\mathrm{x}}_{r}(0)=\mathrm{x}_{r}(0)$ indicates that the system provides the initial GT robot state. See Figure 2 for visualization. Integrating spatial domain RGBD images with their frequency domain counterparts has been shown to enhance image classification performance by accentuating discriminative features [42, 43]. Therefore, we apply the Discrete Cosine Transform [44] to convert RGBD images into the frequency domain. We then concatenate the spatial and frequency domain images as input, enabling $f_{\theta}$ to detect changes in successive, highly-correlated images more effectively than using spatial images alone. Subsequently, KOROL generates the reference trajectory ($\{\hat{\mathrm{x}}_{r}(t)\}_{t=1}^{N}$) by rolling out the dynamics $\mathbf{K}$. We feed these trajectories into the pre-trained inverse dynamic controller [6], which computes the required action $\mathrm{a}(t)$ using $\hat{\mathrm{x}}_{r}(t)$ and $\hat{\mathrm{x}}_{r}(t+1)$.
Updating of Koopman Operator $\mathbf{K}$.
Subsequent updates to the Koopman operator $\mathbf{K}$ are necessitated by changes in the predicted object features $\hat{\mathrm{x}}_{o}(t)$, because $\mathbf{K}$ is optimized for these specific robot states and object features. See pseudocode in Alg. 1 for details. During the training of $f_{\theta}$ from line 7 to line 15, the dynamics $\mathbf{K}$ initially computed at line 3 may no longer be optimal for the new object features, prompting a need for recalculation. However, recalculating the object features across the entire training dataset and updating $\mathbf{K}$ for every $f_{\theta}$ modification is computationally intensive. Therefore, in KOROL, we defer the updates and recalculate $\mathbf{K}$ every $M$ epoches to balance accuracy with computational efficiency, as detailed from line 16 to line 20.
Algorithm 1 Object Feature Learning and Koopman Operator Updating
1:Require training dataset $D$ with robot states $\mathrm{x}_{r}$ and images $y$, feature extractor $f_{\theta}$, function for calculating Koopman operator $func(\cdot)$
2:$\hat{\mathrm{x}}_{o}\leftarrow f_{\theta}(\mathrm{y})$ for $y$ in $D$ // Predict object features across dataset
3:$\mathbf{K}\leftarrow func(\mathrm{x}_{r},\hat{\mathrm{x}}_{o})$ // Calculate the initial dynamics $\mathbf{K}$
4:for epoch$=1,\ldots,N_{1}$ do
5:     $\tau,t_{0}\sim D$ // Sample the trajectory and beginning time steps
6:     $\mathrm{x}_{r}(t_{0}),y(t_{0})\leftarrow\tau(t_{0})$
7:     $loss\leftarrow 0$
8:     for $i=0,\ldots,N$ do
9:         if $i=0$ then
10:              $\hat{\mathrm{x}}_{r}(t_{0})\leftarrow\mathrm{x}_{r}(t_{0})$; $\hat{\mathrm{x}}_{o}(t_{0})\leftarrow f_{\theta}(y(t_{0}))$ // Predict object feature with feature extractor
11:         end if
12:         $\hat{\mathrm{x}}_{r}(t_{0}+i+1)\leftarrow\mathbf{K}^{\prime}(\hat{\mathrm{x}}_{r}(t_{0}+i),\hat{\mathrm{x}}_{o}(t_{0}+i))$ // Predict the next states with $\mathbf{K}$
13:         $loss\leftarrow loss+\left\|\mathrm{x}_{r}(t_{0}+i+1)-\hat{\mathrm{x}}_{r}(t_{0}+i+1)\right\|^{2}$ // Calculate and sum the loss using Equation 6
14:     end for
15:     Update the feature extractor $f_{\theta}$ to minimize $loss$
16:     if epoch %$\%\ \ M$ = 0 then
17:         $\hat{\mathrm{x}}_{o}\leftarrow f_{\theta}(\mathrm{y})$ for $y$ in $D$
18:         $\mathbf{K}\leftarrow func(\mathrm{x}_{r},\hat{\mathrm{x}}_{o})$ // Update the Koopman operator
19:     end if
20:end for
Multi-tasking Koopman Operator.
While a robot’s state space remains consistent when using the same robot platform, object state spaces typically vary across different tasks. For example, a prior Koopman manipulation study [6] includes a 15-DoF tool use task and a 7-DoF door opening task. Due to the differences in object state space definition and dimensions, Koopman operators trained for different tasks can not be shared, limiting their scalability. In KOROL, we propose training object features $\hat{\mathrm{x}}_{o}(t)$ to serve as a universal interface for representing length-varied object states across tasks, enabling the generalization of $\mathbf{K}$ to multiple tasks. Moreover, these object features act as latent conditional vectors that differentiate among tasks. Thus, as long as the feature extractor can identify useful object features, it is possible to use datasets from various tasks to train a single multi-task Koopman operator $\mathbf{K}_{\mathrm{multi}}$.